OpenStudy (anonymous):

Help please!!! P(x)=x^2011+x^1783-3x^1707+2x^341+3x^2-3. Find the remainder when you divide P(x) by x^3-x.

OpenStudy (anonymous):

3x^2 + x - 3 This took quite a while. Easier to do with binomial theorem. But still :p

OpenStudy (anonymous):

Is there a faster way to do this with out the dividing? Surely there is a pattern...

OpenStudy (anonymous):

\[presentable\ answer:\ \sum_{0}^{1004}x^{2n}\]

OpenStudy (amoodarya):

P(x)=x^2011+x^1783-3x^1707+2x^341+3x^2-3 =(x^3-x)Q(x)+R for finding R put x^3=x

OpenStudy (amoodarya):

another way to answer is solving system of equation first find remainder for x second for x-1 third for x+1 now P(x)=x^2011+x^1783-3x^1707+2x^341+3x^2-3=x(x-1)(x+1)q(x)+ax^2+bx+c first put x=0 2nd put x=1 3rd put x=-1 to find a,b,c

OpenStudy (ranga):

^^^^ the quickest way to find the remainder.

OpenStudy (amoodarya):

\[x^{2011}+x^{1783}-3x^{1707}+2x^{341}+3x^{2}-3=xq(x)+r1\rightarrow put \ x=0 \\-3=r1\] \[x^{2011}+x^{1783}-3x^{1707}+2x^{341}+3x^{2}-3=(x-1)k(x)+r2\rightarrow put \ x=1 \\ 1+1-3+2+3-3=r2\rightarrow r2=1\] \[x^{2011}+x^{1783}-3x^{1707}+2x^{341}+3x^{2}-3=(x+1)g(x)+r3\rightarrow put \ x=-1 \\ -1-1+3-2+3-3=r3\rightarrow r3=-1\] \[x^{2011}+x^{1783}-3x^{1707}+2x^{341}+3x^{2}-3=(x-1)x(x+1)h(x)+ax^2+bx+c\rightarrow \\put \ x=0 \\ a(0)+b(0)+c=-3\\ put \ x=1\\a(1)+b(1)+c=1\\ put \ x=-1\\a(+1)+b(-1)+c=-1\\so find a,b,c\]

OpenStudy (anonymous):

ok this is very good thank you. but... if you could explain in baby terms lol that would be great. why did you put (x-1)x(x+1)h(x)+ax^2+bx+c??

OpenStudy (amoodarya):

when you divide to degree one , remain is in degree 0 when you divide to degree 2 , remain is in degree 1 like ax+b when you divide to degree 3 , remain is in degree 2 like this ax^2+bx+c

ganeshie8 (ganeshie8):

beautiful !

OpenStudy (anonymous):

ok then before that why is it (x-1)x(x+1)??

OpenStudy (ranga):

You are asked to divide by x^3 - x = x(x^2 - 1) = x(x+1)(x-1)

OpenStudy (anonymous):

ohhhhh silly me

OpenStudy (ranga):

When you divide a polynomial by a third degree polynomial you will get a quotient and a remainder. The remainder will be one degree less than the denominator and hence ax^2 + bx + c.

OpenStudy (ranga):

P(x) / x(x+1)(x-1). Let q(x) be the quotient and R(x) the remainder. Then, P(x) = x(x+1)(x-1) * Q(x) + R(x) P(0) = 0 + R(0) P(1) = 0 + R(1) P(-1) = 0 + R(-1)

OpenStudy (anonymous):

ok how do i get a, b, and c from there?

OpenStudy (ranga):

R(x) = ax^2 + bx + c R(0) = c R(1) = a + b + c R(-1) = a - b + c P(0) = -3 P(1) = 1 + 1 - 3 + 2 + 3 - 3 = 1 P(-1) = -1 - 1 + 3 - 2 + 3 - 3 = -1 P(0) = R(0) so c = -3 P(1) = R(1) so a + b + c = 1 P(-1) = R(-1) so a - b + c = -1

OpenStudy (ranga):

c = -3 a + b - 3 = 1 or a + b = 4 a - b -3 = -1 or a - b = 2 add the equations: 2a = 6; a = 3 3 + b = 4 b = 1 R(x) = 3x^2 + x - 3

OpenStudy (anonymous):

thanks a bunch!!!!!!

OpenStudy (ranga):

You are welcome. Click on the Best Response for @amoodarya